OpenStudy (misty1212):

Find the complex number that satisfies \(|z-2i|=2,arg(z+2)<\frac{\pi}{4}\)

OpenStudy (misty1212):

my solution: \[\arg(z+2)=\frac{\pi}{4}\implies z+2=a+ai\implies z=a-2+ai\] \[|z-2i|=2\implies (a-2)^2+(a-2)^2=4\implies a=2+\sqrt2\] answer \[z=2+(2+\sqrt2)i\] @phi

OpenStudy (phi):

yes , that's good

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (anonymous):

@phi @misty1212 what section would this fall under for me too review? Complex numbers?